Find the equations of the lines passing through the point(4,-3) and cutting off intercepts on axes whose sum is 6.

Grade:11

1 Answers

Arun
6 years ago
Let 
Equation of line
x/a + y/b = 1
a +b =6
Line passes through (4,-3)
4/a - 3/b = 1
4b - 3a = ab
4(6-a) - 3a = a(6-a)
24 - 4a - 3a = 6a -a²
a²- 13a +24 = 0
Find a from here and then b
Now you can find the equation easily